Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Intern. region
Vertex Tracker
Drift chamber
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of Friday Dec 18 Background Remediation meeting

Tom Mattison gave an overview of the BABAR Protection system. Compared to earlier incarnations, its main feature is to allow to distinguish problems during stored beams and injection. It worked fine during the last machine run. Tim Meyer showed the SVTRAD module that will be implemented in January and has the same logic as Tom’s box. Another BABAR input is available (for the calorimeter ?). Tom Himel pointed out that all signals particpating to the protection logic should :

a-be visible from MCC

b_issue a warning when reaching 70% of the trip limit

A. McKemey gave a status of both protection and operation aspects of the calorimeter. Plans are in preparation for integration PIN diodes as a protection input to PEP-II . A calorimeter module will be installed near PEP-II in January. The data volume at the input of FX extraction is constant in EMC but the time spent in the ROMs depend on the activity and the threshold will need to be raised in case of high background. Same limitations exist for recontrscutiion since there is little or no discriminating power in the EMC between good photons and photons from the background inside the correct time window. No other strong limitaion was found. Study the physcis impact vs the EMC threshold. Compare the reconstruction threshold with the ROM threshold. Prepare a plan for EMC-PIN diode integration into the protection system

H. Band gave a report on the IFR operation and protection aspects. All chambers inside steel should be sufficiently protected from background. The one close the coil are protected by the calorimeter. The troublesome

Chambers are thus layer 18. There is a coupling in the HV distribution ssytem between layer 17 and 18. Study decoupling options. The operation of the chamber is OK up to 40 kHz/cm**2. Data volume is constant with background. Study TDC output (an OR of channels). FE damage have been seen as a result of spikes for the IFR module near PEP-II. Still investigating.

The other protection talks from SVT, DRC and DCH will be given on Jan 8. Documentation already exists for the latter two.